### v3

# Learning When to Think: Shaping Adaptive Reasoning in R1-Style Models via Multi-Stage RL

Abstract
Large reasoning models (LRMs) are proficient at generating explicit, step-by-step reasoning sequences before producing final answers. However, such detailed reasoning can introduce substantial computational overhead and latency, particularly for simple problems. To address this overthinking problem, we explore how to equip LRMs with adaptive thinking capabilities, enabling them to dynamically decide whether to engage in explicit reasoning based on problem complexity. Building on R1-style distilled models, we observe that inserting a simple ellipsis ("…") into the prompt can stochastically trigger either a thinking or no-thinking mode, revealing a latent controllability in the reasoning behavior. Leveraging this property, we propose AutoThink, a multi-stage reinforcement learning (RL) framework that progressively optimizes reasoning policies via stage-wise reward shaping. AutoThink learns to invoke explicit reasoning only when necessary, while defaulting to succinct responses for simpler tasks. Experiments on five mainstream mathematical benchmarks demonstrate that AutoThink achieves favorable accuracy–efficiency trade-offs compared to recent prompting and RL-based pruning methods. It can be seamlessly integrated into any R1-style model, including both distilled and further fine-tuned variants. Notably, AutoThink improves relative accuracy by 6.4% while reducing token usage by 52% on DeepSeek-R1-Distill-Qwen-1.5B, establishing a scalable and adaptive reasoning paradigm for LRMs.
Project Page: https://github.com/ScienceOne-AI/AutoThink.
1 Introduction
Recently, reasoning-focused Large Language Models (LLMs), also referred to as Large Reasoning Models (LRMs) [40], have demonstrated remarkable progress in solving complex reasoning tasks. Particularly, DeepSeek-R1 [9] uses only outcome-based feedback and incentivizes explicit reasoning capabilities through reinforcement learning (RL) with verifiable rewards. DeepSeek-R1 and its distilled models typically follow the <think> and <answer> format, where the <think> process generates explicit, step-by-step reasoning sequences to support obtaining a final answer during the <answer> phase. We refer to models that follow this Chain of Thought (CoT) [36] prompting scheme as R1-style models. The explicit thinking process, which enables self-reflection, backtracking, and validation, is widely regarded as essential for enhancing reasoning accuracy. Arising from this understanding, a popular paradigm has emerged that improves solution quality by increasing thinking token allocation during inference-time reasoning [17, 39]. However, this paradigm introduces a major bottleneck: excessive thinking token generation leads to high computational cost and latency, raising the overthinking phenomenon, where many reasoning steps are redundant or inefficient [29, 14].
To mitigate overthinking, recent efforts have explored hybrid reasoning and concise reasoning strategies. In the industry, Claude 3.7 Sonnet [2] introduces a controllable reasoning framework that allows the model to switch between standard and extended reasoning modes. Similarly, Qwen3 [30] proposes a thinking control scheme with a "thinking" mode (slow thinking) and a "non-thinking" mode (fast thinking), and provides users with the flexibility to choose whether the model should engage in reasoning or not. In the academic community, parallel research has focused on designing prompt-guided efficient reasoning [41, 24] or training pruning-based models to achieve concise reasoning [12, 6, 42]. While promising, these approaches either rely on manually predefined modes or uniformly prune reasoning steps, which may degrade performance on harder instances. A fundamental question then arises to address the overthinking issue:
Can LLMs learn to adaptively determine thinking fast or slow based on given problems?
Figure 1: Overview of AutoThink Compared to Prior Reasoning Paradigms.
To answer this question, we propose AutoThink, a multi-stage RL framework that enables R1-style LLMs to learn adaptive reasoning behaviors. Unlike prior approaches reliant on hard-coded prompting or external control signals, AutoThink formulates reasoning as a learned dual-mode policy that determines both whether to engage the model’s "thinking" process and how to generate concise reasoning. As illustrated in Figure 1, AutoThink fundamentally differs from manual hybrid prompting and uniform pruning strategies by employing an ellipsis prompt and structured three-stage RL training process that enables adaptive reasoning to emerge. In detail, an ellipsis prompt acts as a controllable entry point for optional reasoning, triggering stochastic switching between thinking and no-thinking modes in R1-style LLMs. Then, the proposed multi-stage RL framework shapes this behavior progressively: Stage 1 stabilizes dual-mode coexistence, Stage 2 reinforces accurate reasoning to enhance solution quality, and Stage 3 prunes redundancy via length-aware rewards. This progression enables the model to allocate reasoning effort adaptively, achieving both accuracy and efficiency. The main contributions are as follows:
  • We identify the ellipsis prompt, a lightweight prompting scheme that activates a stochastic switching behavior in R1-style LLMs between thinking and no-thinking modes.
  • We propose a multi-stage RL framework that trains R1-style LLMs to dynamically modulate their reasoning behaviors according to problem complexity.
  • Experiments on mathematical benchmarks show that AutoThink achieves accuracy–efficiency trade-offs better than existing pruning and compression methods, without sacrificing performance.
2 An Ellipsis Unlocks Random Thinking in R1-Style Models
(a) Accuracy and token usage with standard and ellipsis prompts.
(b) No-Thinking rate by difficulty.
Figure 2: Prompting strategies shape reasoning behavior and computational cost.
2.1 A Surprising Effect of Minimal Prompt Modification
Recent efforts on concise reasoning aim to eliminate unnecessary thought, either via prompting that explicitly bypasses thinking [19], or RL-based training that penalizes long outputs [6]. While effective at shortening responses, these methods enforce uniform brevity regardless of problem complexity. Rather than compressing by default, we pose a subtler question:
Can a small change, perhaps a few tokens, lead R1-style models to decide whether to think?
[…]
Surprisingly, this tiny change leads to a distinct shift in behavior. Without any additional training, the model often generates a closing </think> tag, sometimes immediately, skipping deep thinking entirely, and other times after producing a full derivation. As shown in Figure 2(a), evaluation on Distill-R1-1.5B [9] and DeepScaleR [17] across five mathematical benchmarks shows that ellipsis prompting leads to a modest drop in accuracy, accompanied by a substantial reduction in token usage.
Compared to the no-thinking prompt baseline [19], which suppresses reasoning at the cost of accuracy, the ellipsis prompt triggers a stochastic switch in reasoning mode and provides a more balanced trade-off by preserving reasoning when needed and reducing unnecessary computation.
2.2 Prompting Alone Does Not Enable Difficulty-Aware Thinking
The proposed ellipsis prompt seems to trigger selective reasoning: the model thinks on some inputs but not others. While this behavior appears desirable, it raises a deeper question:
Does the prompt-forcing model choose to engage in deep thinking based on task difficulty?
Ideally, a well-calibrated model should reason more on complex problems and skip unnecessary thinking on simpler ones. To assess this, we divide MATH500 problems into 8 difficulty levels based on the average accuracy of Distill-R1 (standard prompt) over 16 rollouts, with higher accuracy indicating lower difficulty. Figure 2(b) (top) shows the no-thinking rate across these levels. Contrary to expectations, under the ellipsis prompt without additional training, no clear trend emerges—the flat distribution suggests that thinking is unguided and unaffected by problem complexity.
A decreasing no-thinking rate along the difficulty axis reflects a desirable reasoning pattern, in which the model allocates effort based on task difficulty. However, this behavior does not emerge from prompting alone. Even with diverse prompt designs (Appendix A.2), the model failed to exhibit difficulty-aware reasoning. Yet prompt-only control suffers from a core limitation: without feedback, the model lacks a mechanism to learn when the thinking process is needed.
To address this gap, we introduce a multi-stage RL framework that rewards appropriate reasoning behavior and encourages alignment between effort and difficulty. As shown in Figure 2(b) (bottom), the resulting distribution from our final trained model exhibits clear difficulty-aware reasoning.
3 Guiding When to Think via Multi-Stage Reinforcement Learning
We propose AutoThink, a multi-stage RL framework with three training phases that induce difficulty-aware reasoning through progressively refined reward designs. At all stages, we employ the GRPO algorithm with a token-level policy gradient loss [25, 44]. The training objective is:
$\mathcal{J}_{\text{GRPO}}(\theta)=$ $\mathbb{E}_{(q,a)\sim\mathcal{D},\;\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot\mid q)}$ (1)
$\Bigg[\frac{1}{\sum_{i=1}^{G}|o_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\min\Big(r_{i,t}(\theta)\hat{A}_{i,t},\text{clip}\left(r_{i,t}(\theta),1-\varepsilon,1+\varepsilon\right)\hat{A}_{i,t}\Big)\Bigg]$
Here, $o_{i}$ denotes the $i$-th sampled output for a given query $q$; $G$ is the number of sampled outputs per query; $r_{i,t}(\theta)$ is the token-level importance weight, defined as the ratio between the new and old token probabilities; and $\hat{A}_{i,t}$ represents the estimated token-level advantage. The overall loss is normalized by the total number of tokens across all sampled trajectories. A visual overview of the reward mechanisms across the three training stages is illustrated in Figure 1. In the following subsections, we detail the reward design for each stage.
3.1 Stage 1: Preventing Mode Collapse by Batch Reward Balance
To promote efficient reasoning, higher rewards are assigned to correct answers without thinking, and stronger penalties to incorrect ones. Define $\text{think}_{i}\in\{0,1\}$ as an indicator of whether the $i$-th output involves thinking, and $\text{correct}_{i}\in\{0,1\}$ as an indicator of whether it yields the correct answer. Based on these variables, the naive reward assignment is:
$r_{i}^{\text{naive}}=\begin{cases}+1,&\text{if }\text{think}_{i}=1\land\text{correct}_{i}=1,\\ \ \ \ 0,&\text{if }\text{think}_{i}=1\land\text{correct}_{i}=0,\\ +2,&\text{if }\text{think}_{i}=0\land\text{correct}_{i}=1,\\ -1,&\text{if }\text{think}_{i}=0\land\text{correct}_{i}=0.\end{cases}$ (2)
While this reward structure encourages difficulty-aware behavior, it causes instability during early training. The model may collapse into a degenerate policy, either always thinking or always skipping, depending on which yields a higher expected reward in the short term. This limits exploration and hinders later optimization. To mitigate this, we introduce batch-level reward balancing:
Figure 3: Effect of $z$ on $r_{i}^{\text{adj}}$.
Let $z\in[0,1]$ denote the proportion of thinking trajectories in a training batch, and 1z1-z the no-thinking proportion. A target balance ratio $\gamma\in(0,1)$ and penalty slope $\lambda\geq 0$ control the strength of adjustment. For thinking and no-thinking samples, we compute soft penalty factors:
$\delta_{\text{think}}$ $=\min\left(1,\max(0,(z-\gamma)\cdot\lambda)\right),$ (3)
$\delta_{\text{nothink}}$ $=\min\left(1,\max(0,(1-z-\gamma)\cdot\lambda)\right).$ (4)
Each sample $i$ is first assigned an original reward $r_{i}^{\text{naive}}\in\{+2,+1,0,-1\}$ based on its thinking flag and correctness. The final adjusted reward is then:
$r_{i}^{\text{adj}}=\begin{cases}(1-\delta_{\text{think}})\cdot r_{i}^{\text{naive}},&\text{if }\text{think}_{i}=1\land\text{correct}_{i}=1,\\ (1-\delta_{\text{think}})\cdot r_{i}^{\text{naive}}+\delta_{\text{think}}\cdot(-1),&\text{if }\text{think}_{i}=1\land\text{correct}_{i}=0,\\ (1-\delta_{\text{nothink}})\cdot r_{i}^{\text{naive}},&\text{if }\text{think}_{i}=0\land\text{correct}_{i}=1,\\ (1-\delta_{\text{nothink}})\cdot r_{i}^{\text{naive}}+\delta_{\text{nothink}}\cdot(-2),&\text{if }\text{think}_{i}=0\land\text{correct}_{i}=0.\end{cases}$ (5)
The adjusted reward $r_{i}^{\text{adj}}$ introduces a soft, piecewise-linear modulation over the naive reward, resembling a hinge-like transformation. Figure 3 illustrates this behavior under a typical setting with $\gamma=0.5$ and $\lambda=2.0$. When thinking dominates ($z>\gamma$), the reward for thinking samples is softly reduced, especially for incorrect answers. Conversely, when no-thinking is overrepresented ($z\ll\gamma$), no-thinking rewards are suppressed. In both cases, the model is gently pushed to restore balance by favoring the less frequent behavior.
3.2 Stage 2: Reinforcing Reliable Behavior within Dual Modes
After establishing behavioral stability across thinking and no-thinking modes, the second stage focuses on improving task performance within each mode. Specifically, the objective is to enhance reasoning quality when invoked, and to promote accurate responses in the absence of thinking.
To allow the model to refine its behavior without external constraints, we remove the batch-level balancing used in the previous stage and allow free evolution of the reasoning policy. The reward is set directly to the naive definition:
$r_{i}^{\text{adj}}=r_{i}^{\text{naive}}.$ (6)
In this stage, we allocate a larger context budget during training, enabling longer responses when needed. Owing to the regularization established in Stage 1, the proportion of thinking in Stage 2 remains balanced, fluctuating naturally rather than collapsing.
3.3 Stage 3: Pruning Unnecessary Reasoning Paths via Length-Aware Reward
While the relaxed setup in Stage 2 improves accuracy, it also leads to overly long responses. Building on the stability established in prior stages, we now aim to improve reasoning efficiency.
Inspired by GRPO-LEAD[48], we introduce a length-aware reward modulation, encouraging brevity in no-thinking mode and rewarding elaboration only when warranted. Specifically, the adjusted reward in this stage is defined as:
$r_{i}^{\text{adj}}=\begin{cases}r_{i}^{\text{naive}}+\left(-1+e^{-\alpha y_{i}}\right),&\text{if }\text{correct}_{i}=1,\\ r_{i}^{\text{naive}}+\left(1-e^{-\beta y_{i}}\right),&\text{if }\text{correct}_{i}=0.\end{cases}$ (7)
where $y_{i}=\frac{L_{i}-\mu_{q}}{\sigma_{q}}$ is the standardized length of response $i$ within its query group $q$. Here, $L_{i}$ denotes the response length, while $\mu_{q}$ and $\sigma_{q}$ are the group-specific mean and standard deviation of lengths, computed separately for correct and incorrect sample groups. And $\alpha$ and $\beta$ are hyperparameters that control the sensitivity of the shaping term.
Figure 4: Effect of $\alpha,\beta$ on $r_{i}^{\text{adj}}$.
The reward decays with length for correct responses and grows for incorrect ones, encouraging concise success and thorough failure analysis, as an example illustrated in Figure 4. This final stage allows the model to adaptively regulate its reasoning depth, producing succinct responses without significantly compromising reliability.
4 Experiments
4.1 Setup
Datasets and Models
We use the same training data as in DeepScaleR [17], comprising 40K mathematically problems with varying difficulties. Following prior works [47, 33], the evaluation is conducted on five standard math benchmarks: MATH, Minerva, Olympiad, AIME24, and AMC23. We evaluate the applicability of AutoThink on three R1-style models with varying sizes and RL post-training status: DeepSeek-R1-Distill-Qwen-1.5B/7B (abbreviated as Distill-R1-1.5B/7B), and DeepScaleR-Preview-1.5B [17] (abbreviated as DeepScaleR), the state-of-the-art 1.5B reasoning model obtained from Distill-R1-1.5B via context-extended RL at a training budget of up to \$5,000.
Baselines
We benchmark our approach against two classes of baselines designed to promote efficient reasoning. (1) Prompt-only baselines: we apply standard [9], no-thinking [19], and ellipsis (ours) prompting strategies on the base models, following the description illustrated in Figure 2(a). (2) RL-trained baselines: including Concise-RL [6], ShorterBetter [42], and ThinkPrune [12], all of which aim to shorten reasoning traces by RL, but do not explicitly account for adaptive reasoning behavior. Among these methods, only ThinkPrune provides publicly available model checkpoints; we evaluate its two representative variants, iter-2K and 4K. For Concise-RL and ShorterBetter, results are reported as published in their respective papers. (3) Additionally, we include a set of open-source RL-finetuned models based on Distill-R1-1.5B/7B as reference, including Open-RS3-1.5B [5], Still-3-1.5B [22], FastCuRL-1.5B [28], Light-R1-DS-7B [37], AReaL-boba-RL-7B [21], and QwQ-32B [31]. These models are not explicitly optimized for concise reasoning and differ significantly in both training objectives and computational budgets. We report their results for contextual reference only, aiming to highlight differences in design philosophy rather than to draw direct performance comparisons.
Training and Evaluation
All experiments are implemented using the verl framework [27], with most training hyperparameters retained at the default values. For all models, we set the batch size and training context length to $(128,8\text{K})$ in Stage 1, $(64,16\text{K})$ in Stage 2, and $(64,24\text{K})$ in Stage 3. We save model checkpoints at empirically selected steps based on observed convergence throughout the procedure: 220/440/130 for Distill-R1-1.5B, 110/240/60 for DeepScaleR, and 220/450/20 for Distill-R1-7B across Stages. During evaluation, all models use a 32K context window. We sample 16 rollouts per instance with temperature 0.6 and report the average pass@1 accuracy. Reward shaping hyperparameters are set to $\gamma=0.5$, $\lambda=2.0$ for Stage 1, and $\alpha=\beta=0.05$ for Stage 3.
Table 1: (Main Results) Accuracy, Token Usage, and Efficiency Comparison Across Methods.
Method Accuracy (%) Token Usage E-F1(%)
MATH Minerva Olympiad AIME24 AMC23 AVG MATH Minerva Olympiad AIME24 AMC23 AVG AVG
Open-Source R1-Style Model: Train From DeepSeek-R1-Distill-Qwen-1.5B/7B or Even Larger
Open-RS3-1.5B 83.0 26.3 43.3 30.6 63.0 49.2 5578 7579 11626 16651 11052 10497 /
Still-3-1.5B 84.9 28.4 45.0 31.0 64.6 50.8 4208 6021 9470 13399 8788 8377 /
FastCuRL-1.5B 87.9 30.9 49.8 40.8 72.3 56.7 3829 5849 7077 10300 6699 6571 /
Light-R1-DS-7B 92.1 37.6 58.1 62.3 82.4 66.5 3774 5434 8362 12064 7317 7390 /
AReaL-boba-RL-7B 93.4 37.7 62.1 65.4 85.7 68.9 4947 8290 10096 12905 8432 8934 /
QwQ-32B 95.1 45.3 69.0 76.7 95.5 76.3 5547 8650 9445 13970 4222 8367 /
Base Model: DeepSeek-R1-Distill-Qwen-1.5B
Standard Prompt 83.1 26.0 43.7 27.5 62.5 48.6 5622 7688 11555 17322 10981 10633 /
No-Thinking Prompt 70.4 19.1 33.1 15.8 49.0 37.5 1256 628 2426 5793 2535 2528 /
Ellipsis Prompt 78.2 21.9 38.6 25.2 57.2 44.2 4194 4336 7752 13006 7980 7453 /
Concise-RL 81.0 / / 30.0 / / 1965 / / 6752 / / /
ShorterBetter / 27.6 38.4 20.0 56.6 / / 1147 1814 2703 1946 / /
ThinkPrune-iter-2k 82.6 28.1 43.6 26.7 64.9 49.2 1927 2126 3683 5806 3300 3368 9.9
ThinkPrune-4k 83.5 28.4 43.4 28.3 65.4 49.8 2723 3375 5504 8072 5040 4943 18.7
AutoThink-Stage1 79.4 21.4 40.5 27.7 59.0 45.6 3107 3867 7212 11673 6467 6465 0.0
AutoThink-Stage2 85.2 27.2 46.4 31.8 66.6 51.4 3702 5481 8030 12117 7415 7295 31.6
AutoThink-Stage3 84.0 28.1 44.8 34.6 67.0 51.7 2195 3212 5559 9514 5059 5108 39.6
Base Model: DeepScaleR-Preview-1.5B
Standard Prompt 87.6 30.7 50.0 42.3 72.8 56.7 3171 4948 5967 9326 5675 5817 /
No-Thinking Prompt 78.1 21.8 40.9 23.8 58.4 44.6 1285 1217 2461 4682 2372 2403 /
Ellipsis Prompt 85.9 28.9 48.1 42.1 72.0 55.4 2890 4748 5416 9408 5095 5511 /
ThinkPrune-iter-2k 86.3 30.7 48.3 38.7 72.2 55.4 1838 2414 3254 5328 3166 3200 0.0
ThinkPrune-4k 86.5 30.6 48.5 36.5 71.8 54.8 2221 3039 4061 6624 3868 3963 0.0
AutoThink-Stage1 82.1 27.0 45.6 33.5 66.0 50.8 2473 5372 7328 12716 5440 $66$ 0.0
AutoThink-Stage2 87.6 31.8 50.1 42.9 73.9 57.3 2762 4315 5521 8567 5222 5277 7.5
AutoThink-Stage3 85.1 30.5 49.0 41.9 71.9 55.7 1897 3834 5005 9033 4696 4893 0.0
Base Model: DeepSeek-R1-Distill-Qwen-7B
Standard Prompt 92.3 37.6 56.4 52.7 82.8 64.4 3928 5155 8815 13563 7613 7815 /
No-Thinking Prompt 78.2 22.1 40.2 22.7 53.7 43.4 722 486 1434 3269 1433 1496 /
Ellipsis Prompt 91.8 37.6 56.5 51.3 80.9 63.6 3752 4778 8643 13532 7616 7564 /
Concise-RL 90.3 / / 51.7 / / 2041 / / 6632 / / /
ShorterBetter / 44.1 50.7 53.3 75.9 / / 1341 3410 5288 2580 / /
AutoThink-Stage1 89.3 31.8 53.8 52.7 78.2 61.2 1763 $17$ 4798 8515 4397 4274 0.0
AutoThink-Stage2 92.2 38.5 56.2 57.1 83.7 65.5 2519 2980 5797 8676 4925 4979 7.2
AutoThink-Stage3 91.2 38.2 56.4 54.8 83.3 64.8 2146 2838 5498 8051 4645 4635 3.2
4.2 Main Results
Table 1 reports average accuracy and token usage across five mathematical benchmarks. To jointly evaluate reasoning accuracy and efficiency, we introduce the Efficiency-F1 score (E-F1), defined as:
$\text{E-F1}=\left(2\cdot\frac{\Delta_{\text{acc}}\cdot\Delta_{\text{len}}}{\Delta_{\text{acc}}+\Delta_{\text{len}}}\right)\ \text{ if }\text{acc}>\text{acc}_{\text{std}}\text{ and }\text{len}<\text{len}_{\text{std}};\ \text{ else }0$
where the normalized accuracy gain and token reduction are given by:
$\Delta_{\text{acc}}=\frac{\text{acc}-\text{acc}_{\text{std}}}{\text{acc}_{\text{std}}-\text{acc}_{\text{no}}},\quad\Delta_{\text{len}}=\frac{\text{len}_{\text{std}}-\text{len}}{\text{len}_{\text{std}}-\text{len}_{\text{no}}}$
The subscripts std and no refer to the standard and no-thinking baselines. A non-zero E-F1 indicates that the model improves upon the standard baseline in both accuracy and token usage, capturing the extent to which pruning enhances conciseness without degrading performance.
Despite the strong performance of existing open-source models, their outputs are substantially longer, even reaching twice the length of ours at the same model size, suggesting that their gains stem from verbose reasoning but non-adaptive reasoning. Prompt-based baselines (no-thinking and ellipsis) reduce length at the cost of accuracy. RL-based baselines also shorten outputs, but offer limited improvements on Distill-R1 and in some cases even reduce accuracy on DeepScaleR.
In contrast, AutoThink exhibits a staged progression in both accuracy and efficiency. All three stages are consistently trained with the ellipsis prompt as the base prompting strategy. Stage 1 primarily aims to stabilize the activation of reasoning behavior and has minimal impact on performance. Stage 2 leads to accuracy improvements over the standard prompt across all model backbones, demonstrating effective control over when to reason. Stage 3 introduces length-aware pruning, further reducing token usage while minimizing potential performance degradation. On Distill-R1-1.5B, AutoThink-Stage3 achieves 51.7% accuracy with half the token usage of the standard prompt baseline. Remarkably, even on the heavily optimized DeepScaleR, AutoThink-Stage2 further improves performance by 0.6 over the standard prompt while reducing token usage by an additional 10%. However, Stage 3 leads to a slight accuracy drop, likely because DeepScaleR has already undergone extensive optimization. This suggests that additional pruning may be unnecessary on fully optimized models.
4.3 Ablation Study
(a) Effect of Stage 1 reward design on reasoning.
(b) Effect of Stage 2 in performance boosting.
Figure 5: Prompting strategies shape reasoning behavior and computational cost.
We conduct ablations on the reward design of our multi-stage RL framework on Distill-R1-1.5B to assess the necessity of each stage. The performance gains achieved by Stage 2 and the pruning effect of Stage 3 are already reflected in Table 1, in terms of accuracy and token usage. Here, we focus on two key aspects: (1) the role of batch reward balance in Stage 1, and (2) whether skipping Stage 2 and proceeding directly from Stage 1 to Stage 3 yields comparable performance.
Batch Reward Balance Prevents Mode Collapse
To assess the role of batch-level balancing in Stage 1, we examine its impact on stabilizing dual-mode reasoning behavior. Specifically, we plot the average thinking rate across training steps, as shown in Figure 5(a). Under a naive reward, the model rapidly collapses into a thinking mode. Conversely, applying the length-aware reward (with $\alpha=0.05$, $\beta=0$) in naive reward to encourage brevity leads the model to collapse into a degenerate no-thinking mode. In contrast, the batch reward balance mechanism, by enforcing a target thinking ratio via penalty slope $\lambda$, helps stabilize training and supports the coexistence of thinking and no-thinking behaviors. We observe that response length rises and then falls during training, indicating an increasing share of shorter, no-thinking responses. These observations imply that the model implicitly performs reasoning pruning, akin to concise reasoning.
Pruning Without Reinforcement Limits Performance
We investigate the necessity of Stage 2 by applying Stage 3 directly after Stage 1, skipping the reinforcement phase. As shown in Figure 5(b), the complete training pipeline that includes Stage 2 prior to Stage 3 yields a notable boost in both accuracy and response length, followed by effective pruning with minimal performance degradation. In contrast, bypassing Stage 2 results in stagnant accuracy and an eventual increase in response length after an initial decline. In contrast, skipping Stage 2 leads to stagnant accuracy and a rebound in response length. With comparable response lengths, the variant achieves only 47.6% accuracy across five benchmarks, notably lower than the 51.7% from full training. These observations underscore the importance of Stage 2 in establishing stable and discriminative reasoning behaviors that enable reliable pruning in the subsequent stage.
4.4 In-Depth Behavioral and Efficiency Analysis
Figure 6: Keyword usage of reasoning behaviors across thinking and no-thinking modes.
Figure 7: Accuracy vs. Thinking Rate. The numbers indicate response lengths (tokens).
Lexical Patterns in Two Reasoning Modes
We analyze linguistic differences between thinking and no-thinking responses by quantifying the frequency of reasoning-related verbs (e.g., “Wait”, “Alternatively”, “Check”) per 1,000 tokens, capturing how explicit reasoning is manifested in each mode. Following [12], we categorize these keywords into three functional groups on the MATH500 benchmark: (1) Soliloquize & Thinking, reflecting internal deliberation and self-correction, characteristic of R1-style reasoning; (2) Check & Confirm, indicating procedural verification; and (3) Summary & Calculation, marking final deduction and computational closure. As illustrated in Figure 7, AutoThink training substantially reduces soliloquy-like expressions, particularly under the no-thinking mode, indicating a decline in explicit internal deliberation. In contrast, verification and computation-related terms appear slightly more frequently in the no-thinking setting, suggesting a shift toward focused conclusion and validation rather than step-by-step verbalization.
Correlation Between Task Difficulty and Reasoning Tendency
We investigate the relationship between the reasoning behavior and the inherent difficulty of the tasks. As shown in Figure 7, there exists a positive correlation between the thinking rate and task difficulty. To further quantify this relationship, we compute the average accuracy, thinking rate, and response length across all datasets. Here, accuracy serves as a proxy for dataset difficulty. The results indicate that, on more challenging datasets, models tend to invoke explicit reasoning more frequently and produce longer responses. This demonstrates that stronger models do not rely on explicit reasoning as frequently, yet outperform weaker models, highlighting an emergent ability to reason more selectively and efficiently.
Readability and Accuracy of Dual Reasoning Modes
A common concern in reinforcement fine-tuning is that reward-driven optimization may degrade the fluency or coherence of generated reasoning traces. To assess whether AutoThink introduces such effects, we follow the evaluation setup in [12] and compute the perplexity (PPL) over the <think> span traces using Distill-R1-1.5B. For no-thinking variants, PPL is calculated over the segment following </think>. As shown in Table 4.4, the think mode of AutoThink maintains PPL comparable to standard prompting, while the no-think mode achieves the lowest PPL, reflecting more concise and fluent responses. Overall, all variants remain within acceptable readability ranges. Meanwhile, we analyze accuracy and token usage across reasoning modes. The results are also recorded in Table 4.4, no-thinking responses are shorter and more accurate, suggesting effective handling of simpler problems. Thinking-mode responses are longer with slightly lower accuracy, reflecting allocation to harder cases. These results indicate that AutoThink adaptively adjusts reasoning depth based on task difficulty.
Evaluating AutoThink Under Standard and No-Thinking Prompts
We analyze how the trained model responds to the standard and forced-no-think prompts. The forced-no-think prompt is defined as $\n...\n</think>\n\n$, which builds upon the ellipsis prompt but enforces an immediate termination of the thinking phase. The results of Distill-1.5B-AutoThink are presented in Table 4.4. As expected, the standard prompt induces longer reasoning traces and achieves higher accuracy, while the forced no-think prompt reduces token usage at the cost of slight performance degradation. These findings suggest that AutoThink has learned to internally compress its reasoning when appropriate, while retaining the ability to conditionally invoke reasoning via prompting.
Table 2: PPL, Acc & Token Length.
Response PPL Acc (%) Token
Model: Distill-R1-1.5B
Standard Prompt 1.61 83.1 5622
No-Thinking Prompt 1.87 70.4 1256
AutoThink: Think Part 2.29 56.4 5090
AutoThink: No-Think Part 1.50 90.1 1592
Model: DeepScaleR-1.5B
Standard Prompt 2.19 83.1 5622
No-Thinking Prompt 1.85 70.4 1256
AutoThink: Think Part 2.43 63.9 5065
AutoThink: No-Think Part 1.84 89.2 1387
Table 3: AutoThink Performance on Three Prompts.
Prompt MATH Minerva Olympiad AIME24 AMC23 AVG
Accuracy (%)
Ellipsis 84.0 28.1 44.8 34.6 67.0 51.7
Standard 84.4 28.1 45.4 35.0 67.5 52.1
Forced No-Think 83.7 27.2 44.8 32.7 65.9 50.9
Token Usage
Ellipsis 2195 3212 5559 9514 5059 5108
Standard 2679 3534 5726 9862 5243 5408
Forced No-Think 2127 2877 5143 8668 4795 4722
Reasoning Behavior Profiling
Figure 8: Distribution of Reasoning Behaviors Across Models and Reasoning Modes.
To gain a deeper understanding of how reasoning behaviors evolve, we annotate the generated solutions from each model with high-level problem-solving phases using GPT-4o. As illustrated in Figure 8, Distill-R1-1.5B distributes its reasoning effort across many surface-level activities, such as “reformulating the problem” and “understanding the problem.” In contrast, ThinkPrune slightly shifts focus toward answer-finalization routines, while still exhibiting dispersed reasoning patterns. Notably, AutoThink in Think Mode allocates a larger proportion of steps to core reasoning phases, including “computing or simplifying expressions” and “applying known theorems,” suggesting a more targeted and efficient reasoning trajectory. Meanwhile, in No-Think Mode, AutoThink maintains strong task comprehension and delivers concise outputs, dedicating most steps to problem understanding and direct computation. These findings indicate that AutoThink not only reduces redundant steps, but also adapts its reasoning structure based on the selected mode.
Generality Beyond Mathematical Reasoning
To investigate whether AutoThink generalizes beyond mathematical reasoning, we additionally evaluate our models on three non-mathematical benchmarks: (i) GPQA for scientific multi-hop reasoning, (ii) MMLU for general multi-task language understanding, and (iii) Live-Code-Bench for code generation (20250727 release).
As shown in Table 4.4, AutoThink retains competitive accuracy while reducing token usage, indicating that adaptive reasoning behaviors extend beyond math tasks. Stage 2 even surpasses the baseline in accuracy while halving response length, highlighting the transferability of our approach to diverse domains.
Table 4: Performance of AutoThink on non-math benchmarks. Each cell shows Accuracy (%) / Avg. Length.
GPQA MMLU Live-Code-Bench Avg
Distill-1.5B 35.1 / 10026 49.5 / 2727 25.2 / 13372 36.6 / 8708
AutoThink-Stage1 31.5 / 8889 47.7 / 1190 23.8 / 5653 34.3 / 5244
AutoThink-Stage2 37.1 / 8617 48.8 / 1743 24.2 / 9647 36.7 / 6669
AutoThink-Stage3 35.7 / 5659 48.8 / 1300 24.9 / 9054 36.5 / 5337
Additional Analysis.
We further conduct additional analyses on more base models, hyperparameters, training cost, and case studies. Details are presented in Appendix B due to space limitations.
5 Related Works
RL-based Post-Training for LLMs.
Reinforcement fine-tuning (RFT) has been widely adopted to improve the reasoning ability of LLMs [32, 13, 9, 11, 7, 34]. Recent work on RL for LLMs has focused on improving the efficiency and effectiveness of large-scale RL training. Key techniques decoupling the clipping mechanism and introducing dynamic group sampling [44], mitigating value bias over long sequences [46, 45], difficulty-aware advantage reweighting [48], model ensembling [8] and designing minimal-form credit assignment strategies for rewards [4]. In addition, RFT has been shown to explicitly promote self-verification and self-correction behaviors [26, 18], while also supporting optimization of test-time compute [23]. Multi-stage, context-length-extended RL further amplifies the long-chain reasoning ability of R1-style models [17, 28]. In our work, RL is applied to train R1-style models to adaptively control their reasoning behavior, enabling selective thinking guided by multi-stage reward shaping.
Mitigating Overthinking for LLMs.
While RFT improves performance, it may induce overthinking, causing models to generate overly verbose reasoning with limited benefit [29, 14]. [3] address overthinking in R1-style models by using self-generated short CoT as positive signals in DPO, encouraging concise reasoning. [49] mitigate overthinking by training models to terminate with “I don’t know” on unsolvable problems. Recent studies have shown that inserting pseudo-thinking cues into R1-style prompts [19], or manually controlling reasoning based on problem difficulty [10, 38, 15], can suppress the model’s thinking behavior, but resulting in reduced performance. Other studies approach the problem from different perspectives: supervised fine-tuning (SFT) with short CoT responses [43, 20], incorporate response length–aware rewards in RFT [1, 42, 12, 6, 16], or leverage smaller models guide larger ones toward faster reasoning [15, 35]. Inspired by these findings, we first design a minimal prompt that elicits random thinking behavior, then apply multi-stage RL to guide the model to think adaptively based on task difficulty, without using external signals or teacher models.
6 Conclusion & Limitations
This work explores how R1-style LLMs can learn to reason adaptively. We propose AutoThink, a minimal prompting strategy paired with a multi-stage RL framework that enables task-aware thinking. Through stage-wise reward shaping, the model stabilizes reasoning patterns, reinforces effective behaviors, and prunes unnecessary steps. Experiments show that AutoThink achieves favorable accuracy–efficiency trade-offs, outperforming prompting and RL baselines without compromising performance, offering a scalable and controllable approach to efficient reasoning in LLMs.
While AutoThink demonstrates promising adaptive reasoning capabilities, several limitations remain: (1) Reward Hacking: The model may bypass the separation between thinking and answering by embedding reasoning after the </think> tag. As shown in Figure 7, reasoning-related tokens still appear in no-thinking mode, suggesting incomplete behavioral separation. (2) Uncontrolled Reasoning Budget: AutoThink adaptively decides when to think, but cannot control overall response length. Future work could explore budget-aware CoT generation, as seen in recent systems like Qwen3 [30]. (3) Unfiltered Training Data: We directly use the DeepScaleR dataset without filtering by task difficulty. Though simple data selection has shown utility, our focus lies in training design. Integrating curriculum-based filtering may further improve performance.
Acknowledgments and Disclosure of Funding
This work is supported by the Strategic Priority Research Program of Chinese Academy of Sciences under Grant XDA0480303, Young Scientists Fund of The State Key Laboratory of Multimodal Artificial Intelligence Systems ES2P100112, National Natural Science Foundation of China 62402252 and 62536003.
Appendix A Additional Definition and Prompts
A.1 Definition of No-Thinking
In R1-style models (e.g., DeepSeek-R1), Thinking refers to generating explicit, step-by-step reasoning traces enclosed within <think> $\cdots$ </think>, enabling reflection and backtracking. By contrast, we define the No-Thinking mode [19] as immediately closing the <think> tag without producing any substantive reasoning, e.g., <think> </think> before moving to the final answer. This phenomenon, also referred to as Non-Thinking in the Qwen3 Technical Report [30], often emerges under our ellipsis prompt, which stochastically toggles the model between reasoning and shortcut modes and thus serves as a lightweight control signal for studying adaptive reasoning behaviors.
A.2 Additional Prompt Variants
Beyond the prompt variants introduced in Section 2, we further explore an alternative strategy that explicitly encourages the model to self-select its reasoning behavior. Specifically, we augment the original CoT with Think By Difficulty (TBD) prompt with an additional clause, followed by ellipsis prompt to preserve the optional-thinking behavior. As is shown below, where the red text highlights the added clause:
[…]
Despite appending the TBD prompt to explicitly encourage adaptive thinking, we observe no meaningful emergence of selective thinking behavior. As shown in Figure 9 and Table 5, we plot the no-thinking rate across difficulty levels (on MATH500) and report accuracy and token usage across five benchmarks. Interestingly, the addition of the TBD prompt leads to a slight drop in both accuracy and token consumption. This result suggests that prompting alone without any reinforcement signal is insufficient to reliably induce adaptive thinking behavior in Distill-R1 models.
Figure 9: No-Thinking Rate.
Prompt MATH Minerva Olympiad AIME24 AMC23 AVG
Accuracy (%)
Ellipsis Prompt 78.2 21.9 38.6 25.2 57.2 44.2
+ TBD Prompt 78.0 21.3 37.1 22.5 55.4 42.9
Token Usage
Ellipsis Prompt 4194 4336 7752 13006 7980 7453
+ TBD Prompt 3893 3122 6490 11754 8796 6811
Table 5: Comparison Between Ellipsis and TBD Prompts.
Appendix B Extended Experimental Results
B.1 Additional Results on Skywork-OR1-Math-7B
To further assess the generality of our method, we apply the AutoThink framework to Skywork-OR1-Math-7B [^1] https://huggingface.co/Skywork/Skywork-OR1-Math-7B, a state-of-the-art 7B model that achieves strong performance on mathematical reasoning tasks. Pretrained and fine-tuned with rule-based reinforcement learning on math and code tasks, this model represents one of the strongest 7B-scale math solvers. As shown in Table 6, the Ellipsis Prompt has limited effect on this highly optimized model, inducing only a marginal proportion of no-thinking responses, indicating reduced prompt sensitivity due to its deterministic reasoning policy.
Despite the limited prompt sensitivity, Stage 1 training with batch-level contrastive signals effectively captures and amplifies the model’s latent no-thinking behavior, enabling more balanced reasoning patterns to emerge. Subsequent Stages 2 and 3 progressively refine this behavior. The full AutoThink framework is applied sequentially over three stages, trained for 600, 500, and 30 steps, respectively. Notably, the final stage achieves a nearly 60% reduction in reasoning tokens (from 9053 to 3974), while preserving task accuracy with less than a 2% degradation compared to the standard prompting baseline. These lightweight training phases are sufficient to induce substantial improvements in efficiency, even on strong pretrained models like Skywork-OR1-Math-7B.
Table 6: Accuracy and Token Usage Comparison on Skywork-OR1-Math-7B.
Method Accuracy (%) Token Usage
MATH Minerva Olympiad AIME24 AMC23 AVG MATH Minerva Olympiad AIME24 AMC23 AVG
Base Model: Skywork-OR1-Math-7B
Standard Prompt 94.0 41.2 62.1 67.1 88.3 70.5 4669 7402 10102 14242 8849 9053
No-Thinking Prompt 85.5 26.1 48.1 45.6 68.2 54.7 1033 775 2982 6416 2402 2722
Ellipsis Prompt 94.0 41.8 61.8 69.1 88.0 70.9 4542 7399 10093 13813 8819 8933
AutoThink-Stage1 92.9 36.9 59.3 65.4 86.6 68.2 1894 2177 4616 7068 4074 3966
AutoThink-Stage2 94.0 40.0 62.3 63.1 88.9 69.7 2298 3247 5437 8091 4521 4719
AutoThink-Stage3 93.1 38.8 61.1 62.7 88.2 68.8 1768 2287 4622 7372 3820 3974
We further visualize the training dynamics of Stage 1 in Figure 10, including the proportion of thinking responses, as well as the response length and accuracy stratified by thinking versus no-thinking behaviors. At early stages, almost all responses involve explicit reasoning. However, batch-wise balancing gradually promotes the emergence of no-thinking behavior. A clear modality shift occurs between steps 100 and 200, marked by a sharp increase in no-thinking responses, which directly contributes to the reduction in average response length. To explicitly encourage a balanced distribution between thinking and no-thinking responses throughout training, we set the target balance ratio $\gamma=0.5$ in each of the three stages. Interestingly, while the accuracy of thinking responses slightly decreases during this phase, the overall accuracy continues to improve. This divergence suggests that the model is learning to skip unnecessary reasoning on simpler problems, thereby increasing both efficiency and decision quality through adaptive control over its reasoning mode.
Figure 10: Training Curves of Skywork-OR1-Math-7B on Stage1.
B.2 Additional Prompt Evaluation on Qwen3
We further extend our study to the Qwen3-8B [30] model by applying the proposed ellipsis prompting and adaptive training strategy. Table 7 reports accuracy and average token length across benchmarks. The results show that ellipsis prompting encourages a non-negligible amount of no-thinking behavior; however, this tendency does not align perfectly with task difficulty (e.g., AIME problems are significantly harder than MATH500, yet elicit a lower thinking rate). Together with observations on Skywork-OR1-Math-7B (where ellipsis prompting induced only $\sim$0.5% no-thinking behavior), these findings suggest that the AutoThink strategy can also induce autonomous reasoning behavior in Qwen3, with $\sim$13% occurrence of no-thinking responses under ellipsis prompting.
Table 7: Results of Qwen3-8B with different prompts. Each cell shows Accuracy (%) / Avg. Length.
Qwen3-8B MATH500 Minerva Olympiad AIME24 AMC23 Avg
Standard Prompt 97.0 / 5351 53.3 / 7010 73.5 / 11342 86.7 / 14690 88.1 / 10343 79.7 / 9747
Ellipsis Prompt 96.4 / 5109 49.5 / 5315 70.9 / 9891 68.3 / 13349 88.9 / 9858 74.8 / 8704
No-Thinking Prompt 84.1 / 1104 41.2 / 639 50.8 / 2860 26.3 / 6518 60.1 / 2913 52.5 / 2807
Ellipsis Prompt: Thinking Rate 96.9% 67.5% 89.0% 87.5% 96.2% 87.4%
B.3 Hyperparameter Sensitivity
The three-stage framework is intentionally designed to be modular and interpretable, with each stage serving a distinct and simple role: (i) Stage 1 introduces a batch-wise reward balance to prevent mode collapse between thinking and no-thinking behaviors; (ii) Stage 2 focuses purely on reinforcing accuracy within each mode without additional reward shaping; (iii) Stage 3 adds length-aware shaping to encourage brevity for correct responses and elaboration for incorrect ones.
Among these stages, only Stage 1 and Stage 3 involve reward shaping beyond naive correctness. Even in these cases, the formulations remain straightforward and principled. Specifically, Stage 1 balances the modal ratio using a linear penalty controlled by hyperparameters $\gamma$ and $\lambda$, which are set to simple default values rather than finely tuned. As illustrated in Figure 3, the resulting reward curve naturally exhibits a symmetric form. Stage 3 reuses shaping terms ($\alpha$, $\beta$) inspired by GRPO-LEAD, again without introducing any ad hoc modifications.
To further examine the robustness of these choices, we conduct sensitivity analyses of $\gamma$, $\lambda$, and $\alpha$ during training, with results summarized below.
Stage 1 Parameters ($\gamma$, $\lambda$).
In Stage 1, we balance the modal ratio between thinking and no-thinking behaviors using a linear penalty controlled by $\gamma$ and $\lambda$. These hyperparameters were not carefully tuned but set to commonly used values. To illustrate their effect, Table 8 reports the average thinking rate at steps 100 and 200 during training. Increasing $\gamma$ encourages more thinking trajectories, while larger $\lambda$ enforces stricter adherence to the target balance.
Table 8: Thinking rate (%) at checkpoints under different values of $\gamma$ and $\lambda$.
$\gamma{=}0.5,\lambda{=}2$ $\gamma{=}0.2,\lambda{=}2$ $\gamma{=}0.8,\lambda{=}2$ $\gamma{=}0.5,\lambda{=}1$ $\gamma{=}0.5,\lambda{=}4$
Thinking-Rate@step100 62.4 57.8 99.9 71.4 51.7
Thinking-Rate@step200 54.2 51.4 100.0 61.2 48.3
Stage 3 Parameters ($\alpha$, $\beta$).
In Stage 3, the shaping terms ($\alpha$, $\beta$) control the rate of reward decay/growth with respect to response length. Table 9 shows response length at checkpoints under different $\alpha$ values with $\beta$ fixed at 0.05. Larger $\alpha$ accelerates length decay for correct responses, while smaller $\alpha$ relaxes the penalty.
Table 9: Response length under different $\alpha$ values ($\beta=0.05$).
$\alpha=0.05$ $\alpha=0$ $\alpha=0.1$
Response-Length@step100 4734 6174 3623
Response-Length@step200 4120 6322 2894
Discussion.
These results confirm that the shaping functions behave as intended: $\gamma$$\lambda$ modulate the balance between modes in Stage 1, and $\alpha$$\beta$ regulate the brevity of responses in Stage 3. Importantly, the overall training trends remain consistent with the main results, demonstrating the robustness of AutoThink without extensive hyperparameter tuning. Thus, while the overall pipeline appears multi-stage, each stage was deliberately designed with minimal tuning and clear interpretability. Looking ahead, it may be possible to unify these stages through a more holistic reward formulation, enabling the model to learn adaptive reasoning behavior within a single-stage process. We leave this as a promising direction for future work.
B.4 Training Cost Comparison
We compare the training cost of AutoThink with two baseline methods, normalizing all runs to a unified batch size of 128. The results are shown in Table B.4, AutoThink adopts a 3-stage schedule with increasing context lengths and a total of 500 steps, comparable to that of 540 in ThinkPrune. In contrast, ShorterBetter trains in a single stage. While prior methods reduce context length to achieve compression, AutoThink expands it but prunes through shorter response length in Stage 3, resulting in comparable training cost. On H100 clusters with 4 nodes, training the all stages can be completed within one day for 1.5B models, and 2.5 days for 7B. Table 10: Training Cost of Distill-R1-1.5B. Method Steps (Batch Size=128) Context Length AutoThink $\approx 220+220+60=500$ 8K / 16K / 24K ThinkPrune $\approx 80+180+180=540$ 4K / 3K / 2K ShorterBetter $\approx 300$ 6K
Table B.4 provide estimated GPU-hour costs for all methods using Distill-R1-1.5B as the base model, AutoThink operates within the same order of compute as concise baselines such as ThinkPrune and ShorterBetter, yet achieves notably stronger performance. In contrast, DeepScaleR, which primarily aims to maximize performance, requires more than $3\times$ higher compute due to its longer context length and increased RL iterations.
Table 11: Estimated GPU-hour cost of training Distill-R1-1.5B with different methods on H100.
Method GPU Hours Avg ACC / Length
ThinkPrune-iter-2K $\sim$400 49.2 / 3368
ShorterBetter $\sim$200 44.7 / 1915
Distill-1.5B-AutoThink $\sim$700 51.7 / 5108
DeepScaleR $\sim$2200 56.7 / 5817
DeepScaleR-1.5B-AutoThink $\sim$250 57.3 / 5277
Appendix C Addressing Potential Challenges
While AutoThink demonstrates robust improvements, several open challenges remain in the area of reasoning control. We outline possible directions for addressing these limitations below:
Token-Budget Control.
Token-budget constraints have been partially explored in prior works [1, 42], where budget-aware reward functions penalize excessively long completions. Such formulations can be readily integrated with AutoThink in a plug-and-play manner to enforce global compute budgets and further improve efficiency.
Dataset Noise.
The presence of noise in large-scale reasoning datasets can hinder training efficiency. Prior studies [28] suggest that curriculum learning or filtering samples by correctness or difficulty can improve learning quality. These strategies are orthogonal to our reward design and could be combined with AutoThink to further enhance robustness.
Reward Hacking.
A common issue is reward hacking, where the model continues reasoning after the </think> tag. This can be mitigated by explicitly penalizing reasoning-related patterns outside the <think> span, or by rewarding clean separation between thought and final answer. Both strategies can be incorporated into future iterations of our reward function.
Overall, we view these solutions as complementary and composable with our framework. Future work will explore tighter integration of these mechanisms to provide a more comprehensive solution to reasoning control.
Appendix D Case Study
Figure 11,12,13 presents some examples comparing four prompting strategies. For easy problems, AutoThink produces correct answers without explicit reasoning, reflecting effective fast thinking. For medium problems, it may activate both reasoning modes, with thinking and no-thinking responses potentially coexisting. For hard problems, the model engages in deeper, slower reasoning, demonstrating iterative understanding and self-verification before arriving at the correct solution. These observations demonstrate how AutoThink adapts its reasoning to problem difficulty, balancing efficiency and reliability through dynamic control of reasoning depth.
Figure 11: Easy Case: AutoThink solves the problem via no-thinking mode with few tokens.
Figure 12: Medium Case: AutoThink exhibits both thinking and no-thinking modes on the problem.
Figure 13: Hard Case: AutoThink solves the problem via thinking mode with repeated verification.